# When a Reinforcement Learning Agent Encounters Unknown Unknowns

### v1

@@ 1 Introduction
[…]
Consider a scenario that is completely untouched in the literature of reinforcement learning – an agent might surprisingly find she has reached an unknown state where she is even unaware of. The former US Secretary of Defense Donald Rumsfeld referred to this scenario as an unknown unknown in a news conference press . Unknown unknowns broadly exist in security-critical applications, such as autonomous vehicles van2018autonomous , quantitative finance liu2021finrl and economics chang2020decentralized , when the environment is extremely under-explored and uncertain.
@@
This paper proposes the first mathematical model to ground unknown unknowns in reinforcement learning. Although actions are calculated by maximising value functions $Q$ and $V$ that are defined on the aware domain, some of them could move the agent to unknown unknown states in the unaware domain. Based upon this, this paper designs a novel model, episodic Markov decision process with growing awareness (EMDP-GA), which employs a newly designed noninformative value expansion (NIVE) approach to expand value functions to the newly aware states, when the agent surprisingly reaches there. Figure 1 illustrates the expansion of the aware domain, where $\mathcal{S}$ is the whole state space and $\mathcal{S}_{t}$ is the aware domain after exploring $t$ episodes. The initial values are chosen as a noninformative prior tibshirani1989noninformative , such as the averaged values on the previous aware domain, out of respect for the extreme lack of knowledge in this unknown unknown state. We further adapt upper confidence bound momentum Q-learning (UCBMQ) menard2021ucb to the growing awareness for training the EMDP-GA, termed as UCBMQ-GA. UCBMQ employs a momentum term to correct the bias in Q-learning, but relies on the upper bounds of value functions $Q$ and $V$. We prove that our NIVE can preserve this ‘upper-bound’ property when expanding the value functions, which will be invalidated by multiplying NIVE’s expansion by any constant $d<1$. This confirms the efficiency of NIVE as well as our adaptation made in EMDP-GA and UCBMQ-GA.
Extensive theoretical analysis is conducted for verifying our approach. We analyse the regret over the whole course, which is defined as the cumulative reward loss over the whole learning process against the hidden optimal policy. The regret has an upper bound in order of $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{4}SA+H^{2}S\sqrt{T})$, where $S$ is the size of the state space, $A$ is the size of the action space, $T$ is the number of episodes, and $H$ is the length of an episode. This regret bound is sublinear with respect to $T$, and comparable with the state of the art (SOTA), including upper confidence bound-advantage zhang2020almost , upper confidence bound momentum Q-learning menard2021ucb and monotonic value propagation zhang2024settling , despite our exposure to unknown unknowns, while they are not. The computing complexity and space complexity of employing the UCBMQ-GA approach for training the EMDP-GA model are of order $\mathcal{O}(H(S+A)T)$ and $\mathcal{O}(HS^{2}A)$, respectively, which are comparable with the SOTA menard2021ucb . These results collectively suggest that although unknown unknowns are surprising, they can nonetheless be discovered asymptotically properly with a decent speed and an affordable cost.
@@ 2 Related Works
A large volume of literature has been on developing EMDP over the last decade. Neu et al. study the adversarial stochastic shortest path problem in EMDPs neu2012adversarial . Zimin and Neu investigate EMDPs with a layered structure zimin2013online . Dann and Brunskill study EMDPs from the perspective of probably approximately correct (PAC) learning dann2015sample . EMDP algorithms can be roughly classified into model-based and model-free approaches. Model-based approaches learn the EMDP model prior to or concurrently with policy optimisation and require $\Omega(S^{2}AH)$ memory space to store the value function estimators moerland2023model , in which recent developments have achieved lower regret and burn-in costs azar2017minimax ; dann2019policy ; zanette2019tighter ; efroni2019tight ; zhang2020almost ; zhang2021reinforcement ; zhang2024settling ; li2024breaking . Here, burn-in costs are the minimum sample size needed for an algorithm to operate sample-optimally. In contrast, model-free approaches estimate the optimal value function or optimise the policy directly, without explicit model estimation, and require only $O(SAH)$ memory to store these estimators. Extensive theoretical analyses of model-free approaches have also been conducted zhang2020almost ; menard2021ucb ; mao2021near ; dann2021provably ; tiapkin2022dirichlet ; li2023breaking ; agrawal2024optimistic . Amongst them, a remarkable result is a minimax regret lower bound, $\mathop{\min}\{\sqrt{SAH^{3}T},HT\}$, established by Domingues et al. domingues2021episodic , offering a theoretical benchmark.
@@
Similar terms are seen in the literature of choice theory grant2015preference ; piermont2017introspective ; kochov2018behavioral . Reverse Bayesianism is proposed to model the expanding decision universe as awareness grows karni2013reverse ; karni2015probabilistic ; karni2017awareness . Belief formation and responses to unforeseen events, as predicted by reverse Bayesianism, are experimentally studied becker2020reverse . This framework was then applied in economic theory viero2021intertemporal ; chakravarty2022reverse ; karni2021reverse and epistemology zednik2016bayesian . However, no practical approach has been proposed to train an agent to be able to handle unknown unknowns.
@@ 3 Notations and Preliminaries
Episodic Markov decision process (EMDP))
An EMDP is defined by a tuple $M=(\mathcal{S},\mathcal{A},P,r,H,T)$ fiechter1994efficient ; kaufmann2021adaptive , described below. $\mathcal{S}$ and $\mathcal{A}$ are finite state space and finite action space, respectively; mathematically, they are subsets of Euclidean spaces. If an agent at a state $s\in\mathcal{S}$ takes an action $a\in\mathcal{A}$, it will be transited to a new state $s^{\prime}\in\mathcal{S}$. An EMDP agent takes episodic actions for learning – after an episode, the agent comes back to the initial state $s_{1}$, and starts the next episode. Usually, episodes involve the same number of time steps, termed as horizon $H\in\mathbb{N}$. $T\in\mathbb{N}$ is the number of episodes in a learning process. We can thus denote the episode $t$ by the trajectory $(s_{1}^{t},a_{1}^{t},\cdots,s_{H}^{t},a_{H}^{t},s_{H+1}^{t})$. For any episode, at time step $h\in[H]\triangleq\{1,\cdots,H\}$, a transition function $P_{h}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ defines the probability $P_{h}(s^{\prime}|s,a)$ of transition from state $s$ to state $s^{\prime}$ when taking action $a$. This interaction leads to an immediate reward $r_{h}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. Let $\pi_{h}(a|s)$ denote the probability of taking action $a$ at state $s$. A policy model $\pi=(\pi_{1},\cdots,\pi_{H})\in\Pi$ characterises the agent’s action course. All possible policy models constitute a policy space $\Pi=\left\{\pi=(\pi_{1},\cdots,\pi_{H})|\pi_{h}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]\right\}$. We denote the policy used to get episode $t$ by $\pi^{t}\in\Pi$.
@@
The objective of reinforcement learning is to learn a policy that maximises the cumulative rewards, denoted by $V_{1}^{\pi}(s)$. Shreve and Bertsekas prove that, under certain conditions, there exists an optimal deterministic policy $\pi$ satisfy $V_{h}^{*}(s)\triangleq V_{h}^{\pi^{*}}(s)=\mathop{\max}_{\pi}V_{h}^{\pi}(s),\ Q_{h}^{*}(s,a)\triangleq Q_{h}^{\pi^{*}}(s,a)=\mathop{\max}_{\pi}Q_{h}^{\pi}(s,a)$ shreve1978alternative . This result leads to the Bellman optimality equation:
@@
UCBMQ is designed for training an EMDP model menard2021ucb . Its objective is to approximate the optimal policy by controlling the regret. It optimises four functions: estimated Q-value function $Q_{h}^{t}(s,a)$, bias-value function $V_{h,s,a}^{t}(s^{\prime})$, upper bound on the optimal Q-value function $\overline{Q}_{h}^{t}(s,a)$, and upper bound on the optimal value function $\overline{V}_{h}^{t}(s)$. In iteration $t$, the agent interacts with the environment to collect episode $t$, updating the visitation counts $n_{h}^{t}(s,a)$ for each state-action pair $(s,a)$. Next, we update the count-dependent parameters, including the bonus term $\beta_{h}^{t}(s,a)$. Finally, we update the four value functions. The core method of updating $Q_{h}^{t}(s,a)$ retains the standard Q-learning update, which is inspired by the Bellman optimality equation, and adds a momentum correction to remove the bias. $V_{h,s,a}^{t}(s^{\prime})$ is updated as a convex combination of the previous bias-value and the optimistic next-step value, $\overline{Q}_{h}^{t}(s,a)$ is updated by adding $\beta_{h}^{t}(s,a)$ to $Q_{h}^{t}(s,a)$ and $\overline{V}_{h}^{t}(s)$ is updated according to Bellman optimality equation. The procedure of UCBMQ is described in Appendix B. UCBMQ can guarantee a regret bound of $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{4}SA)$, where $H$, $S$, $A$ and $T$ are the horizon, the size of the state space, the size of the action space and the number of learning episodes, respectively. This regret bound matches the lower bound $\Omega(\sqrt{SAH^{3}T})$ domingues2021episodic for large enough $T$.
@@ 4.2 Episodic Markov Decision Processes with Growing Awareness
After $t$ episodes, we denote by $U_{t}$ the unknown unknown states. The extended aware domain is thus $\tilde{\mathcal{S}}_{t}=\mathcal{S}_{t}\cup U_{t}$. EMDP usually defines state space and action space independently. Inheriting this convention, in this paper, the agent selects actions from a fixed action space $\mathcal{A}$, despite the fact that the aware domain is growing, even when the agent is placed at new states. Note that the policy $\pi^{t}$, which is used to get episode $t$, is defined on the aware domain after exploring $t-1$ episodes. Thus, the policy $\pi^{t}=(\pi_{1}^{t},\cdots,\pi_{H}^{t})$ is defined on $\tilde{\mathcal{S}}_{t-1}$; i.e., $\pi_{h}^{t}:\tilde{\mathcal{S}}_{t-1}\times\mathcal{A}\rightarrow[0,1]$ for each $h$. Here, $\pi_{h}^{t}(\cdot|s)$ is a distribution over $\mathcal{A}$. Consequently, $\pi^{t}$ effectively constitutes a policy over the entire state space $\mathcal{S}$.
@@
Exploring in an extremely unexplored, uncertain world requires a ‘homeland’. Space exploration needs an earth. We assume that with high probability, unknown unknowns are no more valuable than the average on known states; a very high-value target still exists but with a high risk. Mathematically, with probability at least $1-\frac{\delta}{3}$, the average value of the known states exceeds that of any unknown unknown state:
$\forall s\in\mathcal{S}-\mathcal{S}_{t},Q^{*}_{h}(s,a)\leq\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}Q^{*}_{h}(\hat{s},a),\ V^{*}_{h}(s)\leq\frac{1}{|\mathcal{S}_{t}|}\sum_{\hat{s}\in\mathcal{S}_{t}}V^{*}_{h}(\hat{s}).$
@@ 4.3 Noninformative Prior for Domain Expansion to Newly Aware States
$Q_{h,avg}^{t}(a)=\frac{1}{|\mathcal{S}_{t}|}\sum_{s\in\mathcal{S}_{t}}Q_{h}^{t}(s,a),\ \ \ \overline{V}_{h,avg}^{t}=\frac{1}{|\mathcal{S}_{t}|}\sum_{s\in\mathcal{S}_{t}}\overline{V}_{h}^{t}(s),\ \ \ V_{h,s,a,avg}^{t}=\frac{1}{|\mathcal{S}_{t}|}\sum_{s^{\prime}\in\mathcal{S}_{t}}V_{h,s,a}^{t}(s^{\prime}),$
$V_{h,a,avg}^{t}(s^{\prime})=\frac{1}{|\mathcal{S}_{t}|}\sum_{s\in\mathcal{S}_{t}}V_{h,s,a}^{t}(s^{\prime}),\ \ \ \text{and}\ \ \ V_{h,a,avg}^{t}=\frac{1}{|\mathcal{S}_{t}|^{2}}\sum_{s,s^{\prime}\in\mathcal{S}_{t}}V_{h,s,a}^{t}(s^{\prime}).$
In many reinforcement learning algorithms, such as UCBMQ menard2021ucb used later, upper bounds $\overline{Q}_{h}^{t}(s,a)$ and $\overline{V}_{h}^{t}(s)$ are used to calculate actions. After the surprising encounter, the NIVE expands the estimators of value functions to newly discovered states by assigning them the average values computed over known states. Crucially, the expanded functions $\tilde{\overline{Q}}_{h}^{t}(s,a)$ and $\tilde{\overline{V}}_{h}^{t}(s)$ remain valid upper bounds for $Q^{*}_{h}(s,a)$ and $V_{h}^{*}(s)$, respectively, as the following lemma. The proof is given in Appendix I (implied in Lemma 9).
@@
Further, we may prove the following corollary. The proof is given in Appendix E.

### v2

@@ Abstract
keywords:
unknown unknown , episodic Markov decision process , safe reinforcement learning
[…]
[uoe]organization=University of Edinburgh,country=United Kingdom
[…]
[ustc]organization=University of Science and Technology of China,country=China
@@ 1 Introduction
Reinforcement learning aims to train an agent that takes sequential actions, moving her state in an environment, for maximising her cumulative rewards received from interactions with the environment [30]. A major model in reinforcement learning is Markov decision process (MDP), assuming the state transitions and rewards from actions do not depend on the history given the current state [35]. Further, episodic MDP (EMDP) allows varying rewards from executing the same action at the same state but in different timings [41]. In parallel, value-based algorithms learn value functions $Q$ and $V$ over the space of $(\text{state},\text{action})$-pairs and the state space, respectively, which then deliver a policy model for the agent, taking the actions that maximise $Q$ [30]. The domain of value functions is naturally restricted by the awareness of all possible states – mathematically, they are defined over all aware states, termed as aware domain in this paper. Correspondingly, the exterior of the aware domain is termed as unaware domain.
Consider a scenario that is completely untouched in the literature of reinforcement learning – an agent might surprisingly find she has reached an unknown state where she is even unaware of. The former US Secretary of Defense Donald Rumsfeld referred to this scenario as an unknown unknown in a news conference [27]. Unknown unknowns broadly exist in security-critical applications, such as autonomous vehicles [33], quantitative finance [22] and economics [5], when the environment is extremely under-explored and uncertain.
@@
This paper proposes the first mathematical model to ground unknown unknowns in reinforcement learning. Although actions are calculated by maximising value functions $Q$ and $V$ that are defined on the aware domain, some of them could move the agent to unknown unknown states in the unaware domain. Based upon this, this paper designs a novel model, episodic Markov decision process with growing awareness (EMDP-GA), which employs a newly designed noninformative value expansion (NIVE) approach to expand value functions to the newly aware states, when the agent surprisingly reaches there. Figure 1 illustrates the expansion of the aware domain, where $\mathcal{S}$ is the whole state space and $\mathcal{S}_{t}$ is the aware domain after exploring $t$ episodes. The initial values are chosen as a noninformative prior [32], such as the averaged values on the previous aware domain, out of respect for the extreme lack of knowledge in this unknown unknown state. We further adapt upper confidence bound momentum Q-learning (UCBMQ) [24] to the growing awareness for training the EMDP-GA, termed as UCBMQ-GA. UCBMQ employs a momentum term to correct the bias in Q-learning, but relies on the upper bounds of value functions $Q$ and $V$. We prove that our NIVE can preserve this ‘upper-bound’ property when expanding the value functions, which will be invalidated by multiplying NIVE’s expansion by any constant $d<1$. This confirms the efficiency of NIVE as well as our adaptation made in EMDP-GA and UCBMQ-GA.
Extensive theoretical analysis is conducted for verifying our approach. We analyse the regret over the whole course, which is defined as the cumulative reward loss over the whole learning process against the hidden optimal policy. The regret has an upper bound in order of $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{4}SA+H^{2}S\sqrt{T})$, where $S$ is the size of the state space, $A$ is the size of the action space, $T$ is the number of episodes, and $H$ is the length of an episode. This regret bound is sublinear with respect to $T$, and comparable with the state of the art (SOTA), including upper confidence bound-advantage [40], upper confidence bound momentum Q-learning [24] and monotonic value propagation [38], despite our exposure to unknown unknowns, while they are not. The computing complexity and space complexity of employing the UCBMQ-GA approach for training the EMDP-GA model are of order $\mathcal{O}(H(S+A)T)$ and $\mathcal{O}(HS^{2}A)$, respectively, which are comparable with the SOTA [24]. These results collectively suggest that although unknown unknowns are surprising, they can nonetheless be discovered asymptotically properly with a decent speed and an affordable cost.
@@ 2 Related Works
A large volume of literature has been on developing EMDP over the last decade. Neu et al. study the adversarial stochastic shortest path problem in EMDPs [26]. Zimin and Neu investigate EMDPs with a layered structure [41] . Dann and Brunskill study EMDPs from the perspective of probably approximately correct (PAC) learning [6]. EMDP algorithms can be roughly classified into model-based and model-free approaches. Model-based approaches learn the EMDP model prior to or concurrently with policy optimisation and require $\Omega(S^{2}AH)$ memory space to store the value function estimators [25], in which recent developments have achieved lower regret and burn-in costs [2, 7, 36, 11, 40, 39, 38, 21]. Here, burn-in costs are the minimum sample size needed for an algorithm to operate sample-optimally. In contrast, model-free approaches estimate the optimal value function or optimise the policy directly, without explicit model estimation, and require only $O(SAH)$ memory to store these estimators. Extensive theoretical analyses of model-free approaches have also been conducted [40, 24, 23, 8, 31, 20, 1]. Amongst them, a remarkable result is a minimax regret lower bound, $\mathop{\min}\{\sqrt{SAH^{3}T},HT\}$, established by Domingues et al. [10], offering a theoretical benchmark.
@@
Similar terms are seen in the literature of choice theory [13, 28, 19]. Reverse Bayesianism is proposed to model the expanding decision universe as awareness grows [15, 16, 17]. Belief formation and responses to unforeseen events, as predicted by reverse Bayesianism, are experimentally studied [3]. This framework was then applied in economic theory [34, 4, 14] and epistemology [37]. However, no practical approach has been proposed to train an agent to be able to handle unknown unknowns.
@@ 3 Notations and Preliminaries
Episodic Markov decision process (EMDP)
An EMDP is defined by a tuple $M=(\mathcal{S},\mathcal{A},P,r,H,T)$ [12, 18], described below. $\mathcal{S}$ and $\mathcal{A}$ are finite state space and finite action space, respectively; mathematically, they are subsets of Euclidean spaces. If an agent at a state $s\in\mathcal{S}$ takes an action $a\in\mathcal{A}$, it will be transited to a new state $s^{\prime}\in\mathcal{S}$. An EMDP agent takes episodic actions for learning – after an episode, the agent comes back to the initial state $s_{1}$, and starts the next episode. Usually, episodes involve the same number of time steps, termed as horizon $H\in\mathbb{N}$. $T\in\mathbb{N}$ is the number of episodes in a learning process. We can thus denote the episode $t$ by the trajectory $(s_{1}^{t},a_{1}^{t},\cdots,s_{H}^{t},a_{H}^{t},s_{H+1}^{t})$. For any episode, at time step $h\in[H]\triangleq\{1,\cdots,H\}$, a transition function $P_{h}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ defines the probability $P_{h}(s^{\prime}|s,a)$ of transition from state $s$ to state $s^{\prime}$ when taking action $a$. This interaction leads to an immediate reward $r_{h}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. Let $\pi_{h}(a|s)$ denote the probability of taking action $a$ at state $s$. A policy model $\pi=(\pi_{1},\cdots,\pi_{H})\in\Pi$ characterises the agent’s action course. All possible policy models constitute a policy space $\Pi=\left\{\pi=(\pi_{1},\cdots,\pi_{H})|\pi_{h}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]\right\}$. We denote the policy used to get episode $t$ by $\pi^{t}\in\Pi$.
@@
The objective of reinforcement learning is to learn a policy that maximises the cumulative rewards, denoted by $V_{1}^{\pi}(s)$. Shreve and Bertsekas prove that, under certain conditions, there exists an optimal deterministic policy $\pi$ satisfy $V_{h}^{*}(s)\triangleq V_{h}^{\pi^{*}}(s)=\mathop{\max}_{\pi}V_{h}^{\pi}(s),\ Q_{h}^{*}(s,a)\triangleq Q_{h}^{\pi^{*}}(s,a)=\mathop{\max}_{\pi}Q_{h}^{\pi}(s,a)$ [29]. This result leads to the Bellman optimality equation:
@@
UCBMQ is designed for training an EMDP model [24]. Its objective is to approximate the optimal policy by controlling the regret. It optimises four functions: estimated Q-value function $Q_{h}^{t}(s,a)$, bias-value function $V_{h,s,a}^{t}(s^{\prime})$, upper bound on the optimal Q-value function $\overline{Q}_{h}^{t}(s,a)$, and upper bound on the optimal value function $\overline{V}_{h}^{t}(s)$. In iteration $t$, the agent interacts with the environment to collect episode $t$, updating the visitation counts $n_{h}^{t}(s,a)$ for each state-action pair $(s,a)$. Next, we update the count-dependent parameters, including the bonus term $\beta_{h}^{t}(s,a)$. Finally, we update the four value functions. The core method of updating $Q_{h}^{t}(s,a)$ retains the standard Q-learning update, which is inspired by the Bellman optimality equation, and adds a momentum correction to remove the bias. $V_{h,s,a}^{t}(s^{\prime})$ is updated as a convex combination of the previous bias-value and the optimistic next-step value, $\overline{Q}_{h}^{t}(s,a)$ is updated by adding $\beta_{h}^{t}(s,a)$ to $Q_{h}^{t}(s,a)$ and $\overline{V}_{h}^{t}(s)$ is updated according to Bellman optimality equation. The procedure of UCBMQ is described in B. UCBMQ can guarantee a regret bound of $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{4}SA)$, where $H$, $S$, AA and $T$ are the horizon, the size of the state space, the size of the action space and the number of learning episodes, respectively. This regret bound matches the lower bound $\Omega(\sqrt{SAH^{3}T})$ [10] for large enough $T$.
@@ 4.2 Episodic Markov Decision Processes with Growing Awareness
After $t$ episodes, we denote by $U_{t}$ the unknown unknown states. The extended aware domain is thus $\tilde{\mathcal{S}}_{t}=\mathcal{S}_{t}\cup U_{t}$. EMDP usually defines state space and action space independently. Inheriting this convention, in this paper, the agent selects actions from a fixed action space $\mathcal{A}$, despite the fact that the aware domain is growing, even when the agent is placed at new states. Note that the policy $\pi^{t}$, which is used to get episode $t$, is defined on the aware domain after exploring t1t-1 episodes. Thus, the policy $\pi^{t}=(\pi_{1}^{t},\cdots,\pi_{H}^{t})$ is defined on $\tilde{\mathcal{S}}_{t-1}$; i.e., $\pi_{h}^{t}:\tilde{\mathcal{S}}_{t-1}\times\mathcal{A}\rightarrow[0,1]$ for each $h$. Here, $\pi_{h}^{t}(\cdot|s)$ is a distribution over $\mathcal{A}$. Consequently, $\pi^{t}$ effectively constitutes a policy over the entire state space $\mathcal{S}$.
@@
Exploring in an extremely unexplored, uncertain world requires a ‘homeland’. Space exploration needs an earth. We assume that with high probability, unknown unknowns are no more valuable than the average on known states; a very high-value target still exists but with a high risk. Mathematically, with probability at least $1-\frac{2\delta}{3}$, the average value of the known states exceeds that of any unknown unknown state:
[…]
@@ 4.3 Noninformative Prior for Domain Expansion to Newly Aware States
$Q_{h,avg}^{t}(a)=\frac{1}{|\mathcal{S}_{t}|}\sum_{s\in\mathcal{S}_{t}}Q_{h}^{t}(s,a),\ \ \ \overline{V}_{h,avg}^{t}=\frac{1}{|\mathcal{S}_{t}|}\sum_{s\in\mathcal{S}_{t}}\overline{V}_{h}^{t}(s),$
$V_{h,s,a,avg}^{t}=\frac{1}{|\mathcal{S}_{t}|}\sum_{s^{\prime}\in\mathcal{S}_{t}}V_{h,s,a}^{t}(s^{\prime}),\ \ \ V_{h,a,avg}^{t}(s^{\prime})=\frac{1}{|\mathcal{S}_{t}|}\sum_{s\in\mathcal{S}_{t}}V_{h,s,a}^{t}(s^{\prime}),$
$\text{and}\ \ \ V_{h,a,avg}^{t}=\frac{1}{|\mathcal{S}_{t}|^{2}}\sum_{s,s^{\prime}\in\mathcal{S}_{t}}V_{h,s,a}^{t}(s^{\prime}).$
In many reinforcement learning algorithms, such as UCBMQ [24] used later, upper bounds $\overline{Q}_{h}^{t}(s,a)$ and $\overline{V}_{h}^{t}(s)$ are used to calculate actions. After the surprising encounter, the NIVE expands the estimators of value functions to newly discovered states by assigning them the average values computed over known states. Crucially, the expanded functions $\tilde{\overline{Q}}_{h}^{t}(s,a)$ and $\tilde{\overline{V}}_{h}^{t}(s)$ remain valid upper bounds for $Q^{*}_{h}(s,a)$ and $V_{h}^{*}(s)$, respectively, as the following lemma. The proof is given in H (implied in Lemma 9).
@@
Further, we may prove the following corollary. The proof is given in E.